# Uncovering the role of semantic and acoustic cues in normal and dichotic listening

### v1

@@ Abstract
Despite extensive research, the precise role of acoustic and semantic cues in complex speech perception tasks remains unclear. In this study, we propose a paradigm to understand the encoding of these cues in electroencephalogram (EEG) data, using match-mismatch (MM) classification task. The MM task involves determining whether the stimulus and response correspond to each other or not. We design a multi-modal sequence model, based on long short term memory (LSTM) architecture, to perform the MM task. The model is input with acoustic stimulus (derived from the speech envelope), semantic stimulus (derived from textual representations of the speech content), and neural response (derived from the EEG data). Our experiments are performed on two separate conditions, i) natural passive listening condition and, ii) an auditory attention based dichotic listening condition. Using the MM task as the analysis framework, we observe that - a) speech perception is fragmented based on word boundaries, b) acoustic and semantic cues offer similar levels of MM task performance in natural listening conditions, and c) semantic cues offer significantly improved MM classification over acoustic cues in dichotic listening task. Further, the study provides evidence of right ear advantage in dichotic listening conditions.
@@
EEG, Match-mismatch classification, multi-modal modeling, Speech perception, acoustics and semantics.I Introduction
@@
Humans speech perception is a complex cognitive process involving the integration of acoustic and contextual cues. The speech envelope is the most commonly used acoustic feature for analyzing neural entrainment and is shown to be synchronized with neural oscillations in the auditory cortex [1]. Furthermore, recent research [2] suggests that neural entrainment to speech is not limited to the speech envelope as it also extends to other acoustic features such as phonemic content. However, a systematic understanding of the speech attributes that are encoded in the brain is somewhat lacking. In this paper, we investigate the importance of word boundary information and relative role of semantics and acoustics in parsing the input speech under normal and dichotic listening conditions.
@@
Continuous natural speech stimuli evoke a brain response that can be detected with neuroimaging techniques like electroencephalography (EEG) or magnetoencephalography (MEG). Electroencephalography (EEG) recording of brain responses is one of the simplest approaches to understand the neural processes involved in speech comprehension. The EEG is a non-invasive neural imaging technique that measures electrical activity in the brain by placing electrodes on the scalp [3]. Multiple studies have demonstrated that EEG, recorded during speech listening tasks, contains information about the speech stimuli [1, 4, 5]. Liberto et al. [1] demonstrates that EEG reflects categorical processing of phonemes in continuous speech. Additionally, Broderick et al. [4] showed that EEG captures linguistic representations such as semantic dissimilarity. Similarly, Soman et al. [5] revealed that the EEG recorded during the listening state carries language information of the stimulus.
@@
The design of studies intended to decode the role of different features during speech perception has been an important research area in auditory neuroscience. Early studies [6, 7, 8, 9] in this area explored linear models to establish a relationship between continuous natural speech and EEG responses, classified as forward, backward, or hybrid models. These models predict either the EEG from speech stimuli or reconstruct the speech from EEG responses. The correlation between the predicted and the reference signals has been used as a measure of neural tracking in numerous studies [10]. However, due to the non-linear nature of the auditory system [11], linear models may not be sufficient for capturing the stimulus-response interaction during speech perception. Deep neural networks have recently been used to compare and analyze EEG responses for spoken stimuli [12, 13, 14, 15].
The match-mismatch (MM) tasks have been proposed as an alternative task to the previously explored forward/backward models [14, 16]. In this case, the goal is to determine if a segment of brain response (like EEG) is aligned with a segment of spoken stimulus. This approach, measured by MM classification accuracy, provides a simpler and more interpretable metric compared to the traditionally employed correlation metrics. The correlation approaches often suffer from low correlation values (typically around -%), limiting their applicability [10]. Additionally, correlation metrics are more susceptible to the influence of outliers compared to a classification model. On the other hand, the match-mismatch task, framed as a binary classification problem, is considered an easier task for models to learn when compared to the more complex regression tasks performed by forward-backward models.
In most of the previous studies using the MM task, the auditory stimulus of a fixed duration (s) is processed through a series of convolutional and recurrent layers [12, 15, 16]. The recent study by Soman et al. [17] proposed a deep learning model to investigate stimulus-response alignment at the sentence level. This work showed the importance of incorporating word-level segmentation information for auditory stimulus-response modelling. The work improved the MM classification accuracy compared to the previous works [12] during natural speech listening.
@@
In addition to analysing speech perception during natural speech listening, we also aim to analyse how the brain comprehends speech in complex listening conditions, such as in dichotic listening. Such auditory attention decoding (AAD) experiments have been used to investigate auditory processing and attention mechanisms in the brain [6, 18]. Auditory attention decoding refers to the process of decoding or identifying the specific auditory stimulus or sound that an individual is paying attention to when presented with a mixture of auditory stimuli. One such complex listening condition is the dichotic listening, in which participants are presented with different auditory stimuli simultaneously in each ear. This capability of the human brain is termed ”The Cocktail party effect”, which refers to the ability of the human brain to selectively focus attention on a specific auditory stimulus while filtering out other competing sounds in a noisy environment. Our study aims to identify the specific cues to which the subject directs their attention during this task. Additionally, we aim to analyze the influence of word boundary information on auditory attention decoding (AAD).
@@
In this study, we propose the MM framework for auditory stimulus-response (EEG) modeling with a sentence-level model that incorporates word segmentation information [17]. This framework is designed to capture the neural responses in both mono-aural and dichotic listening scenarios. For dichotic listening, we formulate MM task as a means to perform auditory attention decoding (AAD), aiming to identify the channel the subject was attending. To conduct this study, we utilize two distinct speech-EEG datasets [4]: one collected during a natural speech listening task and the other during a dichotic listening task.
@@
Our approach involves employing a deep learning network for stimulus-response modeling, which consists of three separate sub-networks: one for processing the EEG signals, one for processing the acoustic features of the stimulus and the other for processing the semantic features of the stimulus. The EEG sub-network comprises a series of convolutional layers, followed by word boundary-based average pooling, and a recurrent layer that incorporates inter-word context. The stimuli sub-network is configured based on the input features.
The acoustic feature used to represent phonological traits is the speech envelope. Additionally, the semantic feature employed is the word2vec [19], which represents words as vectors in a high-dimensional space based on their contextual usage. We conduct a comparative analysis to evaluate the relative contributions of acoustic and semantic features in both natural and dichotic listening conditions. The results indicate that semantic features exhibit higher accuracy in match-mismatch (MM) classification during dichotic listening, while both features perform similarly in natural listening conditions.
This paper extends the preliminary work [17] in multiple ways. The major contributions of this paper are:
@@
Formulate a paradigm for auditory stimulus-response (EEG) modeling through a match-mismatch (MM) task in natural listening as well as in dichotic listening.
@@
In the mono-aural natural speech listening task, the MM task performance of audio signals is enriched by the induction of word boundaries. Further, the textual information also provides statistically similar MM performance compared to the acoustic features.
@@
In the dichotic speech listening task, the MM task performance of textual features is higher than that of the audio signal, indicating that EEG signals encode higher-level semantic information which may be accessed to separate the acoustic channels and pay selective attention.
@@
For both these tasks, we propose a multi-modal sequence model based on LSTM architecture. The performance of the multi-modal model significantly improves over the individual modality of speech, indicating that the EEG signal jointly encodes the semantic and acoustic content of the stimulus. Further, we also illustrate that the proposal significantly improves over previously proposed models for MM task [12].
For dichotic listening, we show evidence of right ear advantage through the MM classification performance.
Impact of the proposed study
@@
The results of these experiments find applications in robotics [20, 21], brain-computer interface (BCI), and hearing aid applications [22]. Hearing loss and the associated use of hearing aids are on the rise in the general population [23]. The performance of hearing aids is significantly impacted in situations where background noise involves multiple talkers [24], known as the cocktail party problem. Addressing the challenge of assisted listening in multi-talker environments could have broad societal benefits in terms of communication and mental health [25]. AAD leverages knowledge of the listener’s auditory intent (attention) to isolate and enhance the desired audio stream, while suppressing others [26, 27, 28].
@@
Moreover, auditory attention detection algorithms could be integrated with passive Brain-Computer Interface (pBCI) systems for applications in education and interactive music performances [29]. An early attempt in this direction was made by Cho et al. [30], who developed an attention enhancement system for ADHD children using EEG biofeedback in a virtual classroom environment. The AAD method suggested in this work can be used to continuously track the attended auditory source and provide feedback to the user [29].
@@ II-A Dataset
We experiment with two publicly available speech-EEG data sets111https://doi.org/10.5061/dryad.070jc released by Broderick et al. [4]. These scenarios include subjects listening to uninterrupted, natural speech and a situation where the subjects are exposed to the cocktail party effect. In the latter case, the subjects listen to two distinct audio streams simultaneously, each presented in a separate ear. For the subsequent discussion, we refer to the first dataset as the natural speech dataset and the second one as the dichotic (termed as ”cocktail party” in the original dataset) dataset.
In both experiments, the stimuli were presented using Sennheiser HD650 headphones with the software provided by Neurobehavioral Systems (http://www.neurobs.com). The participants were instructed to keep their gaze fixed on a crosshair displayed on the screen throughout each trial and to minimize activities such as eye blinking and other motor movements.
II-A1 Natural Speech Dataset
The natural speech dataset contains electroencephalographic (EEG) data recorded from subjects as they listened to continuous speech. The subjects listened to a professional audio-book narration by a single male speaker of a well-known work of fiction. For each subject, the data consists of trials of similar length, each containing 180s of audio. The trials preserved the chronology of the storyline without repetitions or breaks. The sentence start and end time and the word-level annotation of the speech recordings are provided. The word segmentation is obtained using a speech recognition-based aligner [31]. The EEG data were acquired using -channel BioSemi system at a sampling rate of Hz, while the audio data is played at kHz. Overall, the speech-EEG data amounted to a duration of hours.
@@ II-A2 Dichotic Dataset
The dichotic dataset used in this study comprises EEG recordings obtained from subjects. Each participant listened to a total of trials, each lasting seconds. They were presented with two well-known fictional works, where one story was played to the left ear and the other to the right ear. Different male speakers articulated each story. The participants were divided into groups of and individuals (with one subject excluded). Each group of listeners was instructed to focus their attention solely on the story presented in either the left or right ear for all the trials. Following each trial, the participants were required to answer multiple-choice questions about the stories played in left/right ear, each offering four possible choices. To maintain consistency, the audio streams of each story within a trial were normalized to have the same root mean squared (RMS) intensity. To prevent the unattended story from capturing the participants’ attention during silent periods in the attended story, any silent gap exceeding seconds was truncated.
@@ II-B EEG Pre-processing
The EEG pre-processing pipeline utilized in this study for both datasets was based on the CNSP Workshop 2021 guidelines222https://cnspworkshop.net/resources.html and implemented using the EEGLAB software [32]. First, a low-pass Butterworth filter with a cutoff frequency of Hz was applied to the EEG signal. This filter helps attenuate high-frequency noise and artifacts that are unrelated to the neural activity of interest. Next, a high-pass Butterworth filter with a cutoff frequency of Hz was employed to remove DC offsets or drifts in the signal. Following the filtering steps, the EEG data were down-sampled to Hz.
@@
II-C Match-mismatch classification task for natural listening
Our work uses the MM task as the stimulus-response model to study the neural entrainment. Figure 1 illustrates this paradigm in detail. The classification model is trained to maximize the similarity between the EEG and the “attended” speech segment while minimizing the similarity between the same EEG data and the “un-attended” audio. In this study, the speech segment is considered at a sentence-level. We also compare with prior works [12, 33], which perform this task at the sentence level. In the natural listening condition, the stimulus corresponding to the EEG response segment is the matched speech while speech from another sentence in the same session is chosen as the mismatched speech. Selecting mismatched samples from the same trial makes the classification task challenging enough to encourage the model to learn the stimulus-response relationships. This sampling approach also avoids the chance of memorizing the speech features along with their labels. The details of MM task for dichotic listening setup are given in the following sub-section.
II-D Auditory Attention Decoding as a Match-mismatch task
Figure 1 illustrates this paradigm in detail. The segment of speech that the subject is instructed to attend is considered as the match segment, while the segment of speech played in the other ear is considered as the mismatch segment. This task is more challenging than MM tasks on the natural speech dataset. The behavioral experiment that followed the listening experiment showed that the subjects were able to comprehend and understand the segment they attended to while having difficulty answering questions about the unattended speech. We devise a trial-selection based on behavioral scores (Figure 3), which is described in Sec. II-H.

### v2

@@ Abstract
Speech comprehension is an involuntary task for the healthy human brain, yet the understanding of the mechanisms underlying this brain functionality remains obscure. In this paper, we aim to quantify the role of acoustic and semantic information streams in complex listening conditions. We propose a paradigm to understand the encoding of the speech cues in electroencephalogram (EEG) data, by designing a match-mismatch (MM) classification task. The MM task involves identifying whether the stimulus (speech) and response (EEG) correspond to each other. We build a multimodal deep-learning based sequence model (STEM3 - Speech-Text EEG Match-Mismatch Model), which is input with acoustic stimulus (speech envelope), semantic stimulus (textual representations of speech), and the neural response (EEG data). We perform extensive experiments on two separate conditions, i) natural passive listening and, ii) a dichotic listening requiring auditory attention. Using the MM task as the analysis framework, we observe that - a) speech perception is fragmented based on word boundaries, b) acoustic and semantic cues offer similar levels of MM task performance in natural listening conditions, and c) semantic cues offer significantly improved MM classification over acoustic cues in dichotic listening task. The comparison of the STEM3 with previously proposed MM models shows significant performance improvements for the proposed approach. The analysis and understanding from this study allows the quantification of the roles played by acoustic and semantic cues in diverse listening tasks and in providing further evidences of right-ear advantage in dichotic listening.
@@
EEG, Match-Mismatch Classification, Multimodal Modeling, Auditory Attention, Acoustics and Semantics.I Introduction
@@
Human speech perception is a complex cognitive process involving the integration of acoustic and contextual cues. The speech signal carries information in the low-level acoustic descriptors of the signal like the envelope, energy, pitch, modulation rate, intonation and stress, as well as in high-level semantic descriptors of the spoken content, speaker characteristics and environment/ambiance. The speech envelope is the most commonly used acoustic feature for analyzing neural entrainment and is shown to be synchronized with neural oscillations in the auditory cortex [1]. Furthermore, recent research [2] suggests that neural entrainment to speech is not limited to the speech envelope as it also extends to other acoustic features such as phonemic content. However, a systematic understanding of the speech attributes that are encoded in the brain is still an open ended question.
@@
In this paper, we investigate the relative role of semantics and acoustics in parsing the input speech under normal and dichotic listening conditions. Continuous natural speech stimuli evoke a brain response that can be detected with neuro-imaging techniques like electroencephalography (EEG) or magnetoencephalography (MEG). The EEG is a non-invasive neural imaging technique that measures electrical activity in the brain by placing electrodes on the scalp [3]. Multiple studies have demonstrated that EEG, recorded during speech listening tasks, contains information about the speech stimuli [1, 4, 5]. In the context of encoding higher-level information, Liberto et al. [1] demonstrated that EEG signals elicit categorical processing of phonemes in continuous speech. Additionally, Broderick et al. [4] showed that EEG captures linguistic representations such as semantic dissimilarity. Similarly, Soman et al. [5] reported that the EEG recorded during the listening state carries information about the language of the stimulus.
@@
The design of studies intended to decode the role of different features during speech perception has been an important research area in auditory neuroscience. Early studies [6, 7, 8, 9] in this area explored linear models to establish a relationship between continuous natural speech and EEG responses, classified as forward, backward, or hybrid models. These models predict either the EEG from speech stimuli or reconstruct the speech from EEG responses. The correlation between the predicted and the reference signals was used as a measure of neural tracking in numerous studies [10]. However, due to the non-linear nature of the auditory system [11], linear models may be incapable of capturing the stimulus-response interactions during speech perception. Deep neural networks have therefore been explored to compare and analyze EEG responses for spoken stimuli [12, 13, 14, 15].
The forward/backward models predict one component (stimulus/response) given the other component (response/stimulus). However, the response (EEG) is not fully determined by the input spoken stimulus as the brain responses also encode other body functions. Hence, most of the successful forward-backward models only predict -% of the data variance. On the other hand, the match-mismatch (MM) task is a compelling task that can uncover the underlying neural mechanisms [14, 16]. In the MM task, the goal is to determine if a segment of brain response (like EEG) matches with a segment of the spoken stimulus that evoked that response. Given the binary nature of the MM task, the design of models is somewhat easier compared to the tedious regression tasks undertaken by forward/backward models. The performance in MM task, measured by MM classification accuracy, provides a simpler and more interpretable metric compared to the traditionally employed correlation metrics. Further, the MM task accuracy achieved by the machine-learning models [17] is high, as even a partial encoding of the stimulus enables the models to predict the MM score.
@@
In addition to analyzing speech perception during natural speech listening, which is typically explored in MM tasks, this paper also delves into the analysis of how the brain perceives speech in complex listening conditions, such as in dichotic listening. Auditory attention decoding refers to the process of decoding or identifying the specific auditory stimulus or sound that an individual is paying attention to when presented with a mixture of auditory stimuli. Such auditory attention decoding (AAD) experiments have been used to investigate auditory processing and attention mechanisms in the brain [6, 18]. One such listening condition is dichotic listening, in which participants are presented with different auditory stimuli simultaneously in each ear. This capability of the human brain is termed ”The Cocktail party effect”, which refers to the ability of the human brain to selectively focus attention on a particular auditory stimulus, while filtering out other competing sounds in a noisy environment. Our study aims to identify the specific cues to which the subject directs their attention during this task. Additionally, we aim to analyze the influence of word boundary information on auditory attention decoding (AAD).
I-A Related works
In earlier studies employing the mismatch (MM) task, auditory stimuli of a fixed duration ( seconds) were typically processed using a combination of convolutional and recurrent neural network layers, as demonstrated in works such as [12, 16, 15]. More recent research has advanced this approach by leveraging distinct EEG responses – Thornton et al. [19] utilized speech-evoked frequency-following responses, while Thornton et al. [20] focused on high-frequency gamma band responses, both achieving notable performance improvements. The role of self-supervised speech representations and contextual text embeddings in the MM task was explored by Wang et al. [17], highlighting their significance in enhancing model performance. Meanwhile, Borsdorf et al. [21] demonstrated that multi-head attention mechanisms offer superior classification accuracy for this task. In a distinct contribution, Soman et al. [22] proposed a deep learning model to investigate stimulus-response alignment at the sentence level during natural speech listening. Their study underscored the importance of incorporating word-level segmentation information.
In dichotic listening, prior works [23] had analysed the semantic processing and more recent works [24, 25] focused on interhemispheric connectivity and sex-related differences in selective attention. However, our work analyzes dichotic speech perception by posing it as an MM task, where the match data is constructed with the stimulus-EEG pair that the subject was attending to, while the mismatch data is constructed with the unattended pair. In this manner, our proposed STEM3 framework models the normal and dichotic listening in a unified framework, allowing us to contrast the roles played by acoustic and semantic cues in these two conditions.
I-B Scope and Contributions
@@
In this study, we propose the MM framework for auditory stimulus-response (EEG) modeling with a sentence-level model that incorporates word segmentation information. This framework is designed to capture the neural responses in both mono-aural and dichotic listening scenarios. To conduct this study, we utilize three distinct speech-EEG datasets [4]: two of them collected during a natural speech listening task and the other during a dichotic listening task.
@@
Our approach employs a deep learning network for stimulus-response modeling, which consists of three separate sub-networks: one for processing the EEG signals, one for processing the acoustic features of the stimulus and the other for processing the semantic features of the stimulus. The EEG sub-network comprises a series of convolutional layers, followed by word boundary-based average pooling, and a recurrent/transformer layer that incorporates inter-word context. The proposed model is termed Speech-Text EEG Match-Mismatch Model (STEM3).
The major contributions of this work are:
@@
Formulate a paradigm for auditory stimulus-response (EEG) modeling through the MM task in natural listening as well as in dichotic listening, and provide a unified framework to understand the relative role of acoustic and semantic cues that are encoded in the EEG data.
@@
In the mono-aural natural speech listening task, the MM task performance of audio signals is enriched by the induction of word boundaries. Further, the textual information provides statistically similar MM performance compared to the acoustic features.
@@
In the dichotic speech listening task, the MM task performance of textual features is significantly higher than those from the audio signal, indicating that EEG signals encode higher-level semantic information.
@@
For both these tasks, the performance of the multimodal STEM3 significantly improves over the speech-EEG models, indicating that the EEG signal jointly encodes the semantic and acoustic contents of the stimulus. The comparison with prior works [21, 17] also illustrates that the STEM3 significantly outperforms these models.
I-C Impact of the proposed study
@@
The results of these experiments find applications in robotics [28, 29], brain-computer interface (BCI), and hearing aids [30]. Hearing loss and the associated use of hearing aids are on the rise in the general population [31]. The performance of hearing aids is significantly impacted in situations where the background noise involves multiple talkers [32], known as the cocktail party problem. Addressing the challenge of assisted listening in multi-talker environments could have broad societal benefits in terms of communication and mental health [33]. AAD leverages knowledge of the listener’s auditory intent (attention) to isolate and enhance the desired audio stream, while suppressing others [34, 35, 36].
@@
The auditory attention detection algorithms could be integrated with passive Brain-Computer Interface (pBCI) systems for applications in education and interactive music performances [37]. An early attempt in this direction was made by Cho et al. [38], who developed an attention enhancement system for ADHD children using EEG biofeedback in a virtual classroom environment. The AAD method suggested in this work can be used to continuously track the attended auditory source and provide feedback to the user [37].
@@ II-A Dataset
We experiment with two publicly available speech-EEG datasets111https://doi.org/10.5061/dryad.070jc, released by Broderick et al. [4]. These scenarios include subjects listening to uninterrupted, natural speech and a situation where the subjects participate in a dichotic listening task. In the dichotic case, the subjects listen to two distinct audio streams presented simultaneously, one in each of the two ears. For the subsequent discussion, we refer to the first dataset as the natural speech dataset (small) and the second one as the dichotic dataset. In both experiments, the stimuli were presented using Sennheiser HD650 headphones with the software provided by Neurobehavioral Systems (http://www.neurobs.com). The participants were instructed to keep their gaze fixed on a crosshair displayed on the screen throughout each trial and to minimize activities such as eye blinking and other motor movements.
To further evaluate and validate our model’s performance in natural settings, we experimented with the larger publicly available dataset, “SparrKULee” [39]. SparrKULee consists of natural speech recordings from a larger pool of subjects and in a non-English setting (Native speakers listening to Flemish language recordings). This expanded subject pool enables us to test on subjects, who are not part of the training dataset. We use the training, validation and test splits following the definitions given in [39] and refer to these experiments as natural speech dataset (large).
II-A1 Natural Dataset - Small
The natural speech dataset small contains electroencephalographic (EEG) data recorded from subjects as they listened to continuous speech. The subjects listened to a professional audio-book narration by a single male speaker of a well-known work of fiction. For each subject, the data consists of trials of similar length, each containing s of audio. The trials preserved the chronology of the storyline without repetitions or breaks. The sentence start/end time and the word-level transcription of the speech recordings are provided. The word segmentation is obtained using a speech recognition-based aligner [40]. The EEG data were acquired using -channel BioSemi system, sampled at Hz, while the audio data is sampled at kHz. Overall, the speech-EEG data amounted to a duration of hours.
@@ II-A2 Dichotic Dataset
The dichotic dataset used in this study comprises EEG recordings obtained from subjects. Each participant listened to a total of trials, with each trial lasting seconds. They were presented with two well-known fictional works, where one story was played to the left ear and the other to the right ear. Different male speakers articulated each story. The participants were divided into groups of and individuals (with one subject excluded). Each group of listeners was instructed to focus their attention solely on the story presented in either the left or right ear for all the trials. Following each trial, the participants were required to answer multiple-choice questions about the stories played in left/right ear. To maintain consistency, the audio streams of each story within a trial were normalized to have the same root mean squared (RMS) intensity. To prevent the unattended story from capturing the participants’ attention during silent periods in the attended story, any silent gap exceeding seconds was truncated.
II-A3 Natural Dataset - Large
The SparrKULee dataset [39] is a speech-evoked auditory repository of EEG, comprising channel EEG recordings from adult (- years) participants (all hearing thresholds dB SPL, for - Hz), with Dutch/Flemish as their native language. The subjects were recruited and presented with speech stimuli in Dutch, for a duration ranging between and minutes, divided into to recordings (i.e., an uninterrupted period in which a participant listens to a stimulus), totaling hours of EEG data. The EEG was collected using a BioSemi ActiveTwo system with 64 active Ag-AgCl electrodes, placed according to - electrode system, with a sampling rate of Hz, where as the speech stimuli was sampled at kHz. More details about this data can be found in [39].
@@ II-B EEG Pre-processing
The EEG pre-processing pipeline utilized in this study for both Natural and Dichotic datasets was based on the CNSP Workshop 2021 guidelines222https://cnspworkshop.net/resources.html and implemented using the EEGLAB software [41]. First, a low-pass Butterworth filter with a cutoff frequency of Hz was applied to the EEG signal. This filter helps attenuate high-frequency noise and artifacts that are unrelated to the neural activity of interest. Next, a high-pass Butterworth filter with a cutoff frequency of Hz was employed to remove DC offsets or drifts in the signal. Following the filtering steps, the EEG data were down-sampled to Hz.
@@
For the Natural Dataset - Large, we utilized the pre-processed data available at https://doi.org/10.48804/K3VSND. This data underwent standard pre-processing steps, consistent with those applied to the datasets mentioned earlier, resulting in a final sampling rate of Hz, as detailed in [39].
II-C MM classification task for natural listening
Our work uses the MM task as the stimulus-response model to study the neural entrainment. The classification model is trained to maximize the similarity between the EEG and the “attended” speech segment while minimizing the similarity between the same EEG data and the “un-attended” audio. In this study, the speech segment is considered at a sentence-level. We also compare with prior works [12, 42], which perform this task at the sentence level. In the natural listening condition, the time-aligned stimulus corresponding to the EEG response segment is the matched speech, while speech from another sentence in the same session is chosen as the mismatched speech. Selecting mismatched samples from the same trial makes the classification task challenging enough to encourage the model to learn the stimulus-response relationships. This sampling approach also avoids the chance of memorizing the speech features along with their labels.
II-D Auditory Attention Decoding as a MM task
Figure 1 illustrates this paradigm in detail. The segment of speech that the subject is instructed to attend is considered as the match segment, while the segment of speech played in the other ear is considered as the mismatch segment. This task is more challenging than MM tasks on the natural speech dataset. The behavioral experiment that followed the listening experiment showed that the subjects were able to comprehend and understand the segment they attended to while having difficulty answering questions about the unattended speech. We employ a trial-selection based on behavioral scores (Figure 3), which is described in Sec. II-H.